# Multi-Excitation Projective Simulation with a Many-Body Physics-Inspired Inductive Bias

### v3

@@ IV A Physics-Inspired Inductive Bias
V Learning Scenarios
In this section, we apply our methods numerically to three synthetic environments. The code can be found in our GitHub repository [48]. The first environment is a small toy environment that allows us to understand the basic numerical properties of mePS agents with different many-body inductive biases in a controlled setting. The second environment is an extension of the first with more actions and a mechanism for deception. Furthermore, its reward contains a contribution measuring the success of an attempted deception. The third environment is used to demonstrate chain-of-thought explanations in multi-layered mePS agents. It models a coarse-grained scenario for diagnosing broken computers.
V.1 Invasion Game With Distraction
Figure 4: The defender $D$ must guess which door the attacker $A$ will go to based on a set of symbols shown to them, and block $A$. $D$ is rewarded for success and punished for failure.
@@ V.1 Invasion Game With Distraction
An environment achieving this goal is constructed as follows: Each round, the agent obtains a percept/observation of the form $(v_{1},v_{2},v_{3})\in\{0,1,2\dots,9\}^{\times 3}$. Here, each entry $v_{j}$ corresponds to a value of an observable $j$. For each observable $j$ and each value $v_{j}$, we associate an atomic clip denoted “$\mathrm{obs_{j}:v_{j}}$” in the percept layer of our two-layer agents. Therefore, each percept is coupled in by putting three excitations into the corresponding atomic clips in the percept layer. As an action, the agent has to pick exactly one of two doors. These two choices are represented by atomic clips “$a=0$” and “$a=1$” in the action layer of the agent. The action gets coupled out as soon as there is an excitation on one of the atomic action clips.
@@
[…]
@@
The many-body $h$-values of the $(2,1)$-agent are of the form $h^{(n)}_{(2,1)}(\{obs_{j}:v_{j},obs_{k}:v_{k}\},a)$ for all $j<k$. There are $6002\cdot\begin{pmatrix}3\\ 2\end{pmatrix}\cdot 10\cdot 10=$ 600 many-body $h$-values/trainable parameters for this agent. This agent has exactly the right inductive bias because its many-body $h$-values $h^{(n)}_{(2,1)}(\{obs_{1}:v_{1},obs_{2}:v_{2}\},a)$ exactly encode the information needed for the right action.
The $(3,1)$-agent has many-body $h$-values of the form $h^{(n)}_{(3,1)}(\{obs_{1}:v_{1},obs_{2}:v_{2},obs_{3}:v_{3}\},a)$. These are $20002\times 10^{3}=$ 2000 trainable parameters, significantly more than for the $(2,1)$-agent. Its many-body $h$-values distinguish between different values of the distraction $v_{3}$, so it is reasonable to expect that this agent also trains slower.
All agents use the softmax function with $\beta=1.0$ to convert $h$-values to probabilities, and all $h$-values are initialized to $1.0$.
After each action, the agent obtains a reward of $+1$ for a right answer and a harsh negative reward of $-10$ for a wrong answer. For the $(2,1)$- and $(3,1)$- agents, this practically prevents the transition from being sampled again. This allows us to map the advantage of the $(2,1)$-agent concerning the number of trainable parameters to an advantage in training time over the $(3,1)$-agent. This mechanism does not apply to the $(1,1)$-agent, since it has no transitions that can deterministically choose the right action.
We train over 10000 rounds, each consisting of one percept-action pair, and average rewards over 100 consecutive rounds. Furthermore, we average the learning curves over 10 agents using the same inductive bias but different random number generator seeds. The results are shown in Figure 5, and confirm our expectations: The 1-body agent cannot solve the problem, the 2-body agent learns to solve the problem perfectly and learns the fastest. The 3-body agent also learns to solve the problem, but it learns slower. From the standard deviations (shaded areas), we see that the fluctuations are negligible.
Figure 5: The average reward learning curves in the Invasion Game With Distraction for agents with the inductive biases discussed in section V.1. Each curve is averaged over an ensemble of 10 agents.
V.2 Deceptive Invasion Game
@@
The first and third observables are comprised of the same values from the previous environment, while the second observable can take values in the range 10-13 and the range of values for the defender’s actions now mirrors that of the first observable. This has the effect of inflating the number of trainable parameters used by the agent for each of the many-body cases considered previously. The $(1,1)$-agent now has $(10+10+4)\cdot 10=240$ many-body h-values, while the $(2,1)$-agent has $180010^{3}+10\cdot 4\cdot 10+4\cdot 10\cdot 10=$ 1800, and the $(3,1)$-agent has $40004\times 10^{3}=$ 4000.
The first observable is interpreted as the door announced by the attacker. The reward is determined based on even and odd parity cases of the first two observables. A reward of +2 is given to the defender if they choose the door shown to them by the attacker in the even parity case, while in the odd parity case, a reward of +2 is given if the agent chooses the next door over. If the defender chooses any other door then they receive a harsh negative reward of -10 to effectively deter them from selecting that option during future deliberations. In the odd parity case, if the agent picked the door announced by the attacker, they receive an additional penalty of -1 (i.e. -11 in total), interpreted as the defender being deceived by the attacker. The agent also gets an additional punishment of $-1$ for picking the wrong door in the even case.
@@
Figure 6: The average reward learning curves in the Deceptive Invasion Game for agents with the inductive biases discussed in section V.2. Each curve is averaged over an ensemble of 10 agents.
@@
V.3 Computer Maintenance
@@
PC overheating: 1 CPU: 13 physical damage: 18 replace components: 9
files disappearing: 2 SSD: 14 software damage: 19 install missing software: 10
visible markings on components: 3 MoBo: 15 malware: 20 cooldown computer: 11
unexpected shutdowns: 4 PSU: 16 faulty: 21 run antivirus: 12
slow performance: 5 OS: 17 not connected: 22
@@
The agent must learn to navigate and solve 11 different scenarios in this work, which we constrained to have at most 2 elements per category to better demonstrate the savings accrued from reducing the size of the ECM. An example scenario is detailed in Figure 7. It is immediately clear from this example scenario that the feedback loop between the MoBo and the SSD demands that these components and the associated causes all be treated together if the agent hopes to solve the problem – something the mePS agent is better suited to handle. The only hope that the single-excitation agent has of solving the scenario is to exponentially inflate the size of their ECM so that all possible combinations of the environment variables are distributed in all atomic clips. If each environment variable can take sufficiently many values, the previous strategy will undoubtedly fail since the random walk path in the agent’s ECM that represents the correct solution to a given scenario will have a vanishingly small probability of occurring.
@@
An inductive bias agent will be represented with the following notation: $\{n_{s},[n_{hc},n_{c}],[n_{ac},n_{f}]\}$, which we call an inductive bias configuration. Here, $n_{s}$ is the many-body cutoff on the number of symptoms in the inductive bias agent’s percept layer; $n_{hc}$, the cutoff on the number of components in the hidden layer; $n_{c}$, the cutoff on the number of causes in the hidden layer; $n_{ac}$, the cutoff on the number of components in the action layer; and $n_{f}$, the cutoff on the number of fixes in the action layer. Since we are discarding leftover excitations, and because we have a layered, feed-forward architecture, the inductive bias used here corresponds to Inductive Bias 2DP from Appendix B. It differs from Inductive Bias 2FF only in the fact that it discards passive excitations, i.e. $C_{\mathrm{out}}:=C^{(o)}_{\mathrm{out}}$. We can calculate the number of learning parameters $N_{l}$ for each agent from Equation (10) in Appendix A, which we will use throughout the rest of this section.

### v4

@@ 5 A Physics-Inspired Inductive Bias
6 Learning Scenarios
In this section, we apply our methods numerically to three synthetic environments. The code can be found in our GitHub repository [43]. The first environment is a small toy environment that allows us to understand the basic numerical properties of MEPS agents with different many-body inductive biases in a controlled setting. The second environment is an extension of the first with more actions and a mechanism for deception. Furthermore, its reward contains a contribution measuring the success of an attempted deception. The third environment is used to demonstrate chain-of-thought explanations in multi-layered MEPS agents operating in a real-world-inspired setting. It models a coarse-grained scenario for diagnosing broken computers. In each of the environments the unrestricted MEPS agent corresponds to a standard PS agent, which will serve as a baseline comparison to MEPS along with a standard 2-layer tabular Q-learning agent and a newly designed tabular PS-inspired multi-layer Q-learning agent for further benchmarking. Further details about the tabular PS-inspired multi-layer Q-learning agent will come in Subsection 6.3 . We choose the Q-learning algorithm because we believe it is the fairest and most straightforward comparison with our model in terms of explainability and performance, also noting that Q-learning is not the only method that MEPS is competitive with.
6.1 Invasion Game With Distraction
Figure 4: The defender $D$ must guess which door the attacker $A$ will go to based on a set of symbols shown to them, and block AA. $D$ is rewarded for success and punished for failure.
@@ 6.1 Invasion Game With Distraction
An environment achieving this goal is constructed as follows: Each round, the agent obtains a percept/observation of the form $(v_{1},v_{2},v_{3})\in\{0,1,2\dots,9\}^{\times 3}$. Here, each entry $v_{j}$ corresponds to a value of an observable $j$. For each observable $j$ and each value $v_{j}$, we associate an atomic clip denoted $\mathrm{obs_{j}:v_{j}}$” in the percept layer of our two-layer agents. Therefore, each percept is coupled in by putting three excitations into the corresponding atomic clips in the percept layer. As an action, the agent has to pick exactly one of two doors. These two choices are represented by atomic clips “a=0a=0” and “a=1a=1” in the action layer of the agent. The action gets coupled out as soon as there is an excitation on one of the atomic action clips.
@@
To keep the comparison of the cases for different $i$ as clean and simple as possible, we sample the percepts $(v_{1},v_{2},v_{3})$ uniformly i.i.d., meaning that we do not need the forgetting and glow mechanisms in the update rule (this corresponds to $\gamma=0$ and $\eta=1$, respectively). This reduces the update rule to $h^{(n+1)}_{(i,o)}=h^{(n)}_{(i,o)}+r$, with $r$ the reward. For each percept $(v_{1},v_{2},v_{3})$, there is exactly one right action $a$. This action depends non-trivially on both of the first two observables. We pick the right action to be $a=v_{1}+v_{2}\ \mathrm{mod}\ 2$. The value of the third observable, $v_{3}$, is just a useless distraction.
@@
The many-body $h$-values of the $(2,1)$-agent are of the form $h^{(n)}_{(2,1)}(\{obs_{j}:v_{j},obs_{k}:v_{k}\},a)$ for all $j<k$. There are $2\cdot\begin{pmatrix}3\\ 2\end{pmatrix}\cdot 10\cdot 10=600$ many-body $h$-values/trainable parameters for this agent. This agent has exactly the right inductive bias because its many-body $h$-values $h^{(n)}_{(2,1)}(\{obs_{1}:v_{1},obs_{2}:v_{2}\},a)$ exactly encode the information needed for the right action.
The $(3,1)$-agent, which represents a standard PS agent here, has many-body $h$-values of the form $h^{(n)}_{(3,1)}(\{obs_{1}:v_{1},obs_{2}:v_{2},obs_{3}:v_{3}\},a)$. These are $2\times 10^{3}=2000$ trainable parameters, significantly more than for the $(2,1)$-agent. Its many-body $h$-values distinguish between different values of the distraction $v_{3}$, so it is reasonable to expect that this agent also trains slower.
All MEPS agents use the softmax function with $\beta=1.0$ to convert $h$-values to probabilities, and all $h$-values are initialized to $1.0$. The standard tabular 2-layer Q-learning agent has a discount factor of $\lambda=0$ and a learning rate of $\alpha=1$ to match the analogous settings used in the MEPS agents.
After each action, the agent obtains a reward of +1+1 for a right answer and a harsh negative reward of 10-10 for a wrong answer. For the $(2,1)$- and $(3,1)$- agents, this practically prevents the transition from being sampled again. This allows us to map the advantage of the $(2,1)$-agent concerning the number of trainable parameters to an advantage in training time over the $(3,1)$-agent. This mechanism does not apply to the $(1,1)$-agent, since it has no transitions that can deterministically choose the right action.
We train over 10000 rounds, each consisting of one percept-action pair, and average rewards over 100 consecutive rounds. Furthermore, we average the learning curves over 50 agents using the same inductive bias but different random number generator seeds. The results are shown in Figure 5, and confirm our expectations: The 1-body agent cannot solve the problem, the 2-body agent learns to solve the problem perfectly and learns the fastest. The 3-body agent also learns to solve the problem, but it learns more slowly. From the standard deviations (shaded areas), we see that the fluctuations are negligible. We can also see that the standard tabular 2-layer Q-learning agent performs slightly better than the standard PS agent (3-body MEPS), but worse than the 2-body MEPS with the optimal inductive bias. These results confirm that the capacity of the (2,1)-agent to discard the distraction allows it to learn faster than the standard PS and Q-learning agents, which both assign separate h/Q-values for each percept.
Figure 5: The average reward learning curves in the Invasion Game with Distraction for MEPS agents with the inductive biases discussed in section 6.1 and a standard tabular 2-layer Q-learning agent. Note that, in this environment, the 3-body PS is equivalent to standard PS and is used as the comparison of MEPS to standard PS. Each curve is averaged over an ensemble of 50 agents.
6.2 Deceptive Invasion Game
@@
The first and third observables are comprised of the same values from the previous environment, while the second observable can take values in the range 10-13 and the range of values for the defender’s actions now mirrors that of the first observable. This has the effect of inflating the number of trainable parameters used by the agent for each of the many-body cases considered previously. The $(1,1)$-agent now has $(10+10+4)\cdot 10=240$ many-body h-values, while the $(2,1)$-agent has $10^{3}+10\cdot 4\cdot 10+4\cdot 10\cdot 10=1800$, and the $(3,1)$-agent has $4\times 10^{3}=4000$. Note that the $(3,1)$-agent is again representing a standard PS agent here.
The first observable is interpreted as the door announced by the attacker. The reward is determined based on even and odd parity cases of the first two observables. A reward of +2 is given to the defender if they choose the door shown to them by the attacker in the even parity case, while in the odd parity case, a reward of +2 is given if the agent chooses the next door over. If the defender chooses any other door then they receive a harsh negative reward of -10 to effectively deter them from selecting that option during future deliberations. In the odd parity case, if the agent picked the door announced by the attacker, they receive an additional penalty of -1 (i.e. -11 in total), interpreted as the defender being deceived by the attacker. The agent also gets an additional punishment of 1-1 for picking the wrong door in the even case.
@@
Figure 6: The average reward learning curves in the Deceptive Invasion Game for MEPS agents with the inductive biases discussed in section 6.2 and a standard tabular 2-layer Q-learning agent. Similar as in Figure 5 the 3-body PS is equivalent to standard PS and is used as a baseline comparison to MEPS. Each curve is averaged over an ensemble of 50 agents.
@@
The performance of the standard tabular 2-layer Q-learning agent here is almost identical to that in Subsection 6.1, but now the performance gap between it and the standard PS agent (3-body MEPS) is even smaller. This also means that the 2-body MEPS agent has a much greater advantage over the standard tabular 2-layer Q-learning agent than before, which may be attributed to the increased memory size of the agents in this environment and the increased complexity of the reward function in the environment.
6.3 Computer Maintenance
@@
PC overheating: 1 CPU: 15 physical damage: 20 replace components: 11
files disappearing: 2 SSD: 16 software damage: 21 install missing software: 12
visible markings on components: 3 MoBo: 17 malware: 22 cooldown computer: 13
unexpected shutdowns: 4 PSU: 18 faulty: 23 run antivirus: 14
slow performance: 5 OS: 19 not connected: 24
@@
The agent must learn to navigate and solve 44 different scenarios in this work, which we constrained to have at most 3 elements per category to better demonstrate the savings accrued from reducing the size of the ECM. An example scenario is detailed in Figure 7. It is immediately clear from this example scenario that the feedback loop between the MoBo and the SSD demands that these components and the associated causes all be treated together if the agent hopes to solve the problem – something the MEPS agent is better suited to handle. The only hope that the single-excitation agent has of solving the scenario is to exponentially inflate the size of their ECM so that all possible combinations of the environment variables are distributed in all atomic clips, which is essentially the strategy of the unrestricted MEPS agent. If each environment variable can take sufficiently many values, the previous strategy will undoubtedly fail since the random walk path in the agent’s ECM that represents the correct solution to a given scenario will have a vanishingly small probability of occurring.
@@
An inductive bias agent will be represented with the following notation: $\{n_{s},[n_{hc},n_{c}],[n_{ac},n_{f}]\}$, which we call an inductive bias configuration. Here, $n_{s}$ is the many-body cutoff on the number of symptoms in the inductive bias agent’s percept layer; $n_{hc}$, the cutoff on the number of components in the hidden layer; $n_{c}$, the cutoff on the number of causes in the hidden layer; $n_{ac}$, the cutoff on the number of components in the action layer; and $n_{f}$, the cutoff on the number of fixes in the action layer. Since we are discarding leftover excitations, and because we have a layered, feed-forward architecture, the inductive bias used here corresponds to Inductive Bias 2DP from  B. It differs from Inductive Bias 2FF only in the fact that it discards passive excitations, i.e. $C_{\mathrm{out}}:=C^{(o)}_{\mathrm{out}}$. We can calculate the number of learning parameters $N_{l}$ for each agent from Equation (12) in A, which we will use throughout the rest of this section.
@@
We choose the inductive bias agent with configuration $\{2,[3,2],[3,2]\}$ to train on the Computer Maintenance environment. This configuration represents the agent with the smallest ECM, at $N_{l}=114375$ trainable parameters, necessary to solve all of the scenarios considered in the training set; it is expected that this agent will be able to reach near-optimality in the fewest number of steps. We also use the unrestricted agent (serving as the standard PS agent) and a tabular PS-inspired multi-layer Q-learning agent (details to follow), both with $N_{l}=1429968$ trainable parameters, for comparison against the inductive bias agent; a difference of about 13 times the number of trainable parameters. Since the standard PS (unrestricted) and tabular PS-inspired multi-layer Q-learning agents see the whole space, they are coded differently to take advantage of the extra speed that certain array structures are endowed with. They always choose the full percept/intermediate clip/action configuration at the start of each deliberative phase and also use a layered, feed-forward ECM. We can represent the standard PS (unrestricted) agents’ architecture using the configuration $\{N_{s},[N_{c},N_{ca}],[N_{c},N_{f}]\}$ (defined in A). The standard PS (unrestricted) and tabular PS-inspired multi-layer Q-learning agents are expected to reach the optimal policy but using many more steps than the inductive bias agent as they are forced to consider a multitude of irrelevant transitions. Note that the inductive bias architecture requires much more real elapsed time than for the standard PS (unrestricted) or tabular PS-inspired multi-layer Q-learning agents (since the code requires better python control flow), so any comparisons between them will be made on the level of total steps taken on average to reach the maximum reward.
Figure 7 illustrates the process leading up to the agent receiving a reward from the environment. The reward is decomposed into two separate mechanisms that are each applied to different $h$-values, which we call the hypothesis and plausibility rewards. We do this to avoid washing out the percept information that tends to happen when blindly applying the standard PS update rule to intermediate layers. The splitting of the reward motivates a modification of the standard tabular 2-layer Q-learning agent to the multi-layer setting so that it can be compatible with the computer maintenance environment, and thus a fair comparison can be drawn between it and our model. The architecture of PS provides a good general template for how to accomplish this modification, which is actually suitable for any tabular method provided one takes appropriate measures to address the credit assignment problem. The idea is to define a tabular memory corresponding to each layer and take the existing tabular deliberation rule and apply it to each layer, taking the output of each layer and feeding it into the next one as the input. The tabular update rule is then applied to the tabular memory corresponding to each layer. For the Q-learning algorithm, we define the set of Q-tables at time $t$ as $\{Q^{(t)}(C_{0},C_{1}),Q^{(t)}(C_{1},C_{2}),\ldots,Q^{(t)}(C_{n-1},C_{n})\}$, where $C_{0}$ are the percepts, $C_{n}$ are the actions, and $C_{i}$ are the hidden memory elements corresponding to layer $i$. The deliberation and update rule for an arbitrary layer $i$ are then
$C^{*}_{i+1}$ $=\text{arg}\max_{C_{i+1}}{Q^{(t)}(C^{*}_{i},C_{i+1})}$ (8)
$Q^{(t+1)}(C_{i},C_{i+1})$ $=(1-\alpha_{i})Q^{(t)}(C_{i},C_{i+1})+\alpha_{i}\left(R^{(t)}+\lambda_{i}\max_{C^{\prime}_{i+1}}{Q^{(t)}(C^{\prime}_{i},C^{\prime}_{i+1})}\right)\,$, (9)
respectively. Here $\alpha_{i}$ is the learning rate, $\lambda_{i}$ is the discount factor, $C^{*}_{i}$ is a previously selected memory element, and $C^{\prime}_{i}$ is a memory element corresponding to the deliberation chain that originated from the newly generated percept received after the agent’s most recent action, each for layer $i$. Other tabular methods can be similarly defined but we will focus on the one defined above for the remainder of this section. Note that while the method described above can promote the architecture of tabular methods to be functionally similar to that of MEPS, this does not endow the agent with the same semantics as MEPS. The PS-inspired multi-layer Q-learning agent described above, for example, cannot have its deliberation be thought of as a random walk through a network of clips - the reason being that Q-learning picks the highest Q-value to choose its action. Therefore, MEPS should be thought of as conceptually distinct, with the ability to inspire new methods in other parts of machine learning.
The hypothesis reward is measured based on how close each intermediate layer element is to its corresponding partner in the given scenario; it is applied to the $h$-values between the percept and intermediate layers. If those elements match exactly, a reward of +5 is given, otherwise, a large penalty of -10 is applied. An additional component of the reward penalizes the agent by up to -1 if they pick more elements than there are in the given scenario; an attempt to discourage the agent from selecting the maximum number of elements permitted by their inductive bias (becomes more important for increasing many-body cutoff size).
The plausibility reward, which is applied to the $h$-values between the intermediate layer and the action layer, has three different components to it. The first component has the same structure as the hypothesis reward but the values are quadrupled for the penalty on too many elements and it is for the agent’s chosen elements from the action layer instead of the intermediate layer. The second component checks whether the chosen components from the intermediate layer match those chosen from the action layer; a reward of +1 is given if they match perfectly, +0.25 if all of the action layer elements match but more intermediate layer elements are chosen then action layer elements, and -2 otherwise. This component of the reward can be interpreted as an internal consistency mechanism that encourages the agent to form coherent explanations between hypothesis and action. Note that it does not directly refer to the scenario key and is thus a general mechanism that can aid learning on scenarios not considered in the original training set. The third component of the reward checks whether the agent’s chosen fixes make sense with respect to the underlying causes they identified in their explanation. This is judged based on certain causal relationships between what the causes and fixes each refer to, that are put in by hand. For example, implementing the fix ‘replace components’ would be justified if the cause of the problem was suspected to be ‘physical damage’ or ‘faulty,’ since both indicate problems with the hardware that are only fixable by physically removing them. However, if the suspected cause was simply ‘malware,’ then only selecting ‘replace components’ would not be justified because this is a fix one implements when there are hardware problems, not software. Now, for each possible fix there are specific causes that need to be selected for the agent to receive the associated reward of +0.3 divided by the number of fixes specified in the chosen scenario key $n_{\mathrm{fix}}$; this ensures that the agent can converge to the optimal reward since some scenarios require different numbers of fixes to solve than others. A penalty of up to $4/n_{\mathrm{fix}}$ is given if the proper causes are not selected. Aside from the use of $n_{\mathrm{fix}}$, the third component of the plausibility reward is also independent of the scenario key, adding another general mechanism for augmenting learning beyond the training set.
Finally, if all elements in the agent’s explanation exactly match those in the scenario key, the agent receives a big bonus to both rewards of +15 to amplify the probability of selecting this deliberation path again in future episodes. Once this bonus is triggered it signals an end to the episode and a fresh set of symptoms corresponding to another scenario are then selected. Before receiving this signal and before the agent factors in the reward into their update rule definition 6, the reward is first sent to an external reward shaping function defined by Equation (13) whose purpose is to discourage the agent from taking too many steps within an episode, intending to curtail arbitrarily long episodes; details about this function can be found in A.
All MEPS agents for each of their layers use the softmax function with $\beta=1/2$ to convert $h$-values to probabilities and all $h$-values are initialized to 1. The tabular PS-inspired multi-layer Q-learning agent uses the same values for their hyperparameters from the previous Invasion Game environments for each layer. The results of the training are shown in Figure 8. We can see that the inductive bias agent achieves near-optimal values for both the hypothesis and plausibility rewards using far fewer steps but requiring more episodes than both the standard PS (unrestricted) and tabular PS-inspired multi-layer Q-learning agents do, as seen in the total step number per episode curve, which better illustrates the savings from the inductive bias as the two average reward plots do not show how many total steps have elapsed during training. The total step number per episode curve also shows roughly exponential decay, which is indicative of the agent first trying out scenario configurations randomly and then always picking the right configuration immediately. To quantify the difference in step number to the optimal reward, we calculated the total average steps taken over 600 episodes for all agents: the inductive bias agent takes roughly 16766 steps, while the standard PS (unrestricted) and tabular PS-inspired multi-layer Q-learning agents each take roughly 35817 and 115261 steps, respectively. Together with the complexity-theoretic results from D, we only expect this difference in step number to grow. Notice also how the tabular PS-inspired multi-layer Q-learning agent requires 79444 more steps on average than the standard PS (unrestricted) agent does, and that it doesn’t follow a strict exponential decay curve for the average total step number. We suspect the reason for this is that, because the tabular PS-inspired multi-layer Q-learning agent has a deterministic policy with no exploration, it finds the correct scenario relatively quickly but then has difficulty adapting to new scenarios, hence the initial increase in average step number. Although the standard PS (unrestricted) and tabular PS-inspired multi-layer Q-learning agents do converge to the optimal rewards in fewer episodes, it clearly takes more steps to do so because they both have roughly 13 times the number of trainable parameters as the inductive bias agent has. The fact that we get comparable performance from the inductive bias agent using far fewer parameters justifies the relatively small fluctuations around the optimal value for each of the reward curves, which are expected to decrease as more agents are included in the average.
Figure 8: The average hypothesis and plausibility rewards, and step number per episode learning curves in the Computer Maintenance environment for the agent with inductive bias configuration $\{2,[3,2],[3,2]\}$ (IBC 1), the standard PS (unrestricted) agent, and the tabular PS-inspired multi-layer Q-agent. Each agent is trained for 1000 episodes (600 shown), where the number of steps taken within an episode varies from 1 (optimal) to around 800 (worst). Note that the x-axis of the total step number plot is restricted to 600 episodes so that the details of the early episodes can be better resolved, since the curves beyond this point are essentially flat. Each curve is further averaged over an ensemble of 50 agents. The standard deviation around the curves (shaded areas) is used to represent fluctuations but should not be interpreted as points that the individual agents have necessarily visited.
7 A Quantum Motivation of the Inductive Bias and a Path Towards Quantization
In Section 5, we gave a qualitative, conceptual motivation of our inductive biases. There, we identified that the size of the power set $\mathcal{P}(V)$ can lead to exponential costs, because in general all of its elements might be valid excitation configurations for which we have to define $h$-values.
However, the analogy to quantum many-body physics can be worked out more formally, and this in particular suggests a natural quantization for analog quantum computers. Therefore, in Section 7.1 we will work out this formal analogy, while in Section 7.2 we will provide some basic considerations for an actual quantum implementation.
7.1 Full Quantum Motivation